OpenStudy (anonymous):

4(t + 20) = 2/5 (10t + 200) Please show work

OpenStudy (anonymous):

4t+80=4t+80 4t-4t=80-80 0=0 you cannot find the value of t using this

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

20(t+20)=20t+400 40t+400=20t+400 0=0 question is wrong value of t cannot be find using this

OpenStudy (bakonloverk):

thts true u can't find t using this

OpenStudy (anonymous):

Wait is the top answer correct or the bottom

OpenStudy (anonymous):

both are correct